OpenStudy (anonymous):

3x + 2y = 20 y = x + 4 help

OpenStudy (anonymous):

sove by sub methode

OpenStudy (anonymous):

3x + 2(x + 4) = 20 3x + 2x + 8 = 20 5x = 12 x = 12/5 y = 32/5